OpenStudy (anonymous):

MEDAL AND FAN what is the solution to this proportion? 12/15 = 8/x A. 12 B. 9 C. 10 D. 16

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

\[\frac{ 12 }{ 15 } = \frac{ 8 }{ x }\] Do the cross multiplication single step to get rid of the fractions. \[8*15 = 12*x\]

OpenStudy (danjs):

divide both sides by 12 to get x by itself as x = ...